A wave can travel at a speed 11.1 m/s on a wire that is under a tension of 5.3 N. What tension is required for a wave speed of 35.7 m/s in the same wire?



1
Expert's answer
2021-05-26T15:32:38-0400

Explabations & Calculations


  • Wave speed on a tensed thread is given by

"\qquad\qquad
\begin{aligned}
\small v&=\small \sqrt{\frac{T}{\mu}}
\end{aligned}"

  • Since it is the same thread used in both the cases, linear density — "\small \mu" — remains unchanged.
  • Then the speed appears to depend only on the tension in the thread.

"\qquad\qquad
\begin{aligned}
\small v\propto \sqrt T
\end{aligned}"

  • Therefore, if the required tension for the second case is "\small T_1",

"\qquad\qquad
\begin{aligned}
\small \frac{v_1}{v_0}&=\small \frac{\sqrt {T_1}}{\sqrt{T_0}}\\
\small T_1&=\small T_0\cdot\big(\frac{v_1}{v_0}\big)^2\\
&=\small 5.3\,N\times \big(\frac{35.7\,ms^{-1}}{11.1\,ms^{-1}}\big)^2\\
&=\small \bold{54.8\,N}
\end{aligned}"

  • Notice the increased tension as higher the wave speed expected. The thread should be able to withstand this tension for a successful behavior.